OpenStudy (anonymous):

how can i do this? The following function defines a recursive sequence. f(0) = -3 f(1) = 6 f(n ) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? a. -3, 6, -9, -12, … b. -3, 20, -95, 480, … c. -3, 6, -9, 12, … d. -3, -20, -95, -480, …

OpenStudy (anonymous):

n=2?

OpenStudy (jdoe0001):

have you done geometric sequences?

OpenStudy (anonymous):

not much..

OpenStudy (jdoe0001):

lemme rewrite that in a bit different notation\(\huge{ a_o = -3\qquad \qquad a_1=6\\ \quad \\ a_n = -2\cdot a_{n-1}-a_{n-2}}\)

OpenStudy (jdoe0001):

if you want to find the \(\bf n^{th}\) term, you plug that as your "n", so to match against your choices, you got the 1st one and the 2nd one, get the 3rd, n = 2, and the 4th one, n = 3

OpenStudy (anonymous):

so u would juss add one to each nth term?

OpenStudy (anonymous):

a2=-a3??

OpenStudy (jdoe0001):

hmm well... \(\bf \Large {a_o = -3\qquad \qquad a_1=6\\ \quad \\ a_n = -2\cdot a_{n-1}-a_{n-2}\\ \quad \\ \quad \\ a_2 = -2\cdot a_{2-1}-a_{2-2}\\ \quad \\ a_3 = -2\cdot a_{3-1}-a_{3-2}}\)

OpenStudy (anonymous):

i need to find wat a2 and a3 is?

OpenStudy (jdoe0001):

yes, once you find what \(\bf a_2\) is, you can go and find \(\bf a_3\) \(\large\bf a_2 = -2\cdot a_{2-1}-a_{2-2}\implies a_2=-2\cdot a_1-a_0\)

OpenStudy (jdoe0001):

we know what \(\bf a_1 \ and \ a_0\) are, so we can just plug them in, to get \(\bf a_2\)

OpenStudy (anonymous):

a2=-2x6-3 ?

OpenStudy (jdoe0001):

\(\large\bf {a_2 = -2\cdot a_{2-1}-a_{2-2}\implies a_2=-2\cdot a_1-a_0\\ \quad \\ a_2 = -2\cdot (6)-(-3)}\)

OpenStudy (anonymous):

a2=-9

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

i understand it now o-o. thanks so much : ))!!!!

OpenStudy (jdoe0001):

\(\large \bf {a_3 = -2\cdot a_{3-1}-a_{3-2}\implies a_3=-2\cdot a_2-a_1}\)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

a3=18-6 a3=12 !:D

OpenStudy (jdoe0001):

:)